Write and solve a system of equations for a given model / word problem:
1. A total of $10,000 was invested in two bond mutual funds, a junk bond fund and a government bond fund. The junk fund is risky and yields 11% interest. The safer government bond fund yields only 5%. The year’s total income from the two investments was $740. How much was invested in each fund?

Answer by stanbon(57967)

A total of $10,000 was invested in two bond mutual funds, a junk bond fund and a government bond fund.
j + g = 10000
---------------------

The junk fund is risky and yields 11% interest. The safer government bond fund yields only 5%. The year’s total income from the two investments was $740.
0.11j + 0.05g = 740
------------------------
How much was invested in each fund?
---------------------------
Rearrange for elimination:
5j + 5g = 50000
11j+ 5g = 74000
-----------------------
Subtract 1st from 2nd:
6j = 24000
j = $4000 (amount invested in junk)
Since j+g= 10000, g = $6000
================================
Cheers,
Stan H.